# aoas/sm_enkf

### v2

@@ Abstract
The ensemble Kalman Filter (EnKF), as a fundamental data assimilation approach, has been widely used in many fields of earth science, engineering and beyond. However, several theoretical aspects of the EnKF remain unknown, especially when the state variable is of high dimensional accompanied with high resolution observation and physical models. This paper first proposes several high dimensional EnKF (HD-EnKF) methods that provide consistent estimators for the important forecast error covariance and the Kalman gain matrix. It then studies the theoretical properties of the EnKF under both the fixed and high dimensional state variables, which provides the mean square errors of the analysis states to the underlying oracle states offered by the Kalman filter and gives the much needed insight into the roles played by the forecast error covariance on the accuracy of the EnKF. The accuracy of the data assimilation under the misspecified physical model is also considered. Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed HD-EnKF algorithms perform better than the standard EnKF methods as they provide more robust and accurate assimilated results. The HD-EnKF was applied to assimilate sea temperature in the Northwest Pacific, which showed more accurate out-sample performances than the existing methods.
@@ 1 Introduction
Kalman Filter (Kalman, 1960) is a fundamental approach for forecasting and data assimilation for a partially observed dynamic system. Data assimilation is a statistical procedure that combines information from numerical physical model forecasts with observed data to obtain the best possible description of a dynamical system and its uncertainty (Evensen et al., 2022). The assimilated state variables at a time are used to generate the forecast for the next time step. For chaotic systems sensitive to initial values, data assimilation can significantly improve the accuracy of the forecast by using optimized state variables as initial values (Lorenz, 1963). Although much applied in numerical weather and oceanography prediction, Kalman Filter and its later developments (Ensemble Kalman Filter, three and four-dimensional variational assimilation (3D-Var, 4D-Var; Pailleux, 1990; Courtier et al., 1998)) are the methodological core for scientific algorithms over a range of fields, including hydrology, remote sensing data utilization, carbon data assimilation, robotic control and automatic driving. Data assimilation methods also provide the underlying algorithms and predictive techniques for building digital twins for social and economic systems.
@@
Estimating the forecast error covariance is a key in the EnKF, which is made in the EnKF by the sample covariance of the forecast ensembles. Past works on the EnKF found that the sampling error can lead to underestimation of the forecast error covariance and can eventually cause filter divergence(Anderson and Anderson, 1999; Constantinescu et al., 2007). Filter divergence refers to a decreasing ability to correct the ensemble state towards the observations after certain rounds of assimilation. High resolution data assimilation and forecasting techniques are much needed in modern meteorology and atmospheric systems like global carbon calculation and ocean data assimilation, which leads to high dimensional geophysical models with high dimensional state vectors. Besides, along with advances in remote sensing technology in measuring the earth system, there is a growing availability of high spatial resolution satellite observations (Stroud et al., 2010; Guinehut et al., 2012).
@@
The high dimensional state vectors lead to high dimensional  forecast error covariance and Kalman gain matrices. The high dimensionality is known to cause the non-convergence of the sample forecast error covariance (Bai and Yin, 1993), which is felt in the form of filter divergence and unreliable data assimilation in practice. Geophysicists have proposed two methods to counter the problems. One is the localization method (Furrer et al., 2006; Furrer and Bengtsson, 2007) that sets covariances to zero if the geographical distances are bigger than a threshold; and the other is the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term (Wang and Bishop, 2003; Liang et al., 2012; Constantinescu et al., 2007). However, the statistical properties of these EnKF methods are generally lacking, despite their widespread applications. Katzfuss et al. (2020) proposed Gibbs Ensemble Kalman Filter and Smoother, and particle Ensemble Kalman Filter and Smoother algorithms for high dimensional hierarchical dynamic models from a Bayesian perspective.
This paper proposes several high dimensional EnKF (HD-EnKF) methods based on consistent high dimensional covariance estimators and conducts a theoretical analysis of assimilated state vectors. It is shown that consistent estimation of the forecast error covariance leads to the consistency of the Kalman gain matrix by establishing an inequality that bounds the error of estimation of the Kalman Gain matrix by that of the forecast error covariance. This implies that, even at the high dimensional setting, the mean square errors (MSEs) of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators. Thus, consistent recovery of the underlying analysis states can be attained by the proposed HD-EnKF procedure.
We establish the mean square error bounds and weak convergence of the analysis states from the HD-EnKF to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models. The MSE bounds for the multiple-step forecasting and data assimilation are also provided. Simulation studies suggest that the proposed HD-EnKF algorithm could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method. We apply the proposed HD-EnKF to assimilate sea temperature by combining a physical ocean model with satellite observations and use the ARGO data to validate the quality of the assimilated sea temperature. Both the simulation and the empirical study showed advantages of the HD-EnKF approach.
@@
The paper is structured as follows. Section 2 reviews the Kalman Filter and the EnKF. Section 3 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman gain matrix, and establishes the theoretical properties of the EnKF. The effects of imperfect or misspecified models on the accuracy of the EnKF are provided in Section 4. Results of empirical data assimilation based on two well-known geophysical models are reported in Section 5 and the application to the sea temperature data assimilation is presented in Section 6, followed by a conclusion in Section 7. Details of the proposed HD-EnKF algorithms, additional theoretical and simulation results, and technical details are given in the Supplementary Material (SM).
@@ 2 Review on Kalman and Ensemble Kalman Filters
Assuming $\mathbf{M}_{t}$, $\mathbf{Q}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are known, the Kalman Filter provides a two-step alternative updating framework based on the theory of multivariate normal distribution. In the forecast step, suppose that $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p}\left(\mathbf{x}_{t-1}^{a},\boldsymbol{\Sigma}_{t-1}^{a}\right)$, where $\mathbf{x}_{t-1}^{a}$ is the analysis state vector at the previous time $t-1$ and $\mathbf{y}_{1:t-1}$ denotes the observation $\{\mathbf{y}_{i}\}_{i=1}^{t-1}$ up to time $t-1$, the conditional distribution of $\mathbf{x}_{t}$ given $\mathbf{y}_{1:t-1}$ is
@@
is the variance of the forecast error $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$, also called the forecast error covariance matrix.
In the update (assimilation) step, the goal is to update $\mathbf{x}_{t}^{f}$ according to the new observation $\mathbf{y}_{t}$ and to reduce the difference between the true state $\mathbf{x}_{t}$ and the forecast state $\mathbf{x}_{t}^{f}$. The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is Gaussian, which implies that
$\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right),$ (2.4)
where, by the conditional Gaussian distribution formulae,
@@
However, models in earth science or engineering are often nonlinear, that is the evolution model in (2.1) may not be linear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$, where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2). Besides, for large dynamic systems, the dimension $p$ of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store. To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error covariance is no longer updated as (2.3) but by the sample covariance based on $n$ forecast ensembles. It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing.
The EnKF assumes possibly nonlinear model system with linear observation operator:
$\left\{\begin{array}[]{l}\mathbf{x}_{t}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1}\right)+\mathbf{w}_{t},\quad\quad\ \mathbf{w}_{t}\sim\mathrm{N}_{p}(\mathbf{0},\mathbf{Q}_{t}),\\ \mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\boldsymbol{\varepsilon}_{t},\quad\quad\quad\quad\quad\boldsymbol{\varepsilon}_{t}\sim\mathrm{N}_{q}\left(\mathbf{0},\mathbf{R}_{t}\right),\end{array}\right.$ (2.7)
where $\mathcal{M}_{t}$ is the possibly nonlinear model operator, $\mathbf{H}_{t}$ is a linear observation operator. It further assumes that the forecast error $\boldsymbol{\eta}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$ follows a Gaussian distribution given $\mathbf{y}_{1:t-1}$, that is $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}\mid\mathbf{y}_{1:t-1}\sim\mathrm{N}_{p}(\mathbf{0},\boldsymbol{\Sigma}_{t}^{f})$, where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)$. Here $\mathcal{M}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are assumed to be known for the time being as in the Kalman Filter.
@@
    $\mathbf{x}_{t,j}^{f}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1,j}^{a}\right)+\mathbf{w}_{t,j}\text{ and }\hat{\mathbf{x}}_{t}^{f}={n}^{-1}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{f},\text{ where }\mathbf{w}_{t,j}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right).$ (2.8)
@@
    and obtain an estimate of the Kalman gain matrix $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$.
@@
    Calculate the perturbed residuals $\mathbf{d}_{t,j}=\mathbf{y}_{t}+\boldsymbol{\varepsilon}_{t,j}-{\mathbf{H}}_{t}\mathbf{x}_{t,j}^{f}$ where $\{\boldsymbol{\varepsilon}_{t,j}\}_{j=1}^{n}\stackrel{{\scriptstyle IID}}{{\sim}}{\mathrm{N}}\left(\mathbf{0},\mathbf{R}_{t}\right)$, and let $\mathbf{d}_{t}=n^{-1}\sum_{j=1}^{n}\mathbf{d}_{t,j}$. Compute the perturbed analysis states and their average, the assimilated (analysis) state,
@@
The EnKF does not require the expensive recursions (2.3) and (2.5)-(2.6) used in the Kalman Filter. Instead, it estimates the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance. For the case with non-linear model operator as in (2.7), we can still define the underlying analysis state $\mathbf{x}_{t}^{a}$ by (2.5)-(2.6), where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)=\mathbb{E}(\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}\mid\mathbf{y}_{1:t-1})$ is the forecast state and $\boldsymbol{\Sigma}_{t}^{f}$ is the forecast error covariance defined by $\boldsymbol{\Sigma}_{t}^{f}=\operatorname{var}((\mathbf{x}_{t}-\mathbf{x}_{t}^{f})\mid\mathbf{y}_{1:t-1})$. In fact, $\mathbf{x}_{t}^{a}$ defined by (2.5)-(2.6) is the solution to the minimizing problem (2.11)
$J(\mathbf{x})=\frac{1}{2}\big{\{}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\big{\}}$ (2.11)
with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$. For the ease of reference, and to avoid differentiating between the linear and the non-linear cases, we also call $\mathbf{x}_{t}^{a}$ as the analysis state of the Kalman Filter even in the non-linear model case. The next two sections will establish the convergence of the EnKF analysis state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.

### v3

# High Dimensional Ensemble Kalman Filter

@@ Abstract
The Ensemble Kalman Filter (EnKF), as a fundamental data assimilation approach, has been widely used in many fields of the sciences and engineering. When the state variable is of high dimensional accompanied with high resolution observations of physical models, some key theoretical aspects of the EnKF are open for investigation. This paper proposes several high dimensional EnKF (HD-EnKF) methods equipped with consistent estimators for the important forecast error covariance and Kalman Gain matrices. It then studies the theoretical properties of the EnKF under both fixed and high dimensional state variables, which provides one-step and multiple-step mean square errors of the analysis states to the underlying oracle states offered by the Kalman Filter and gives the much needed insight to the roles played by the forecast error covariance on the accuracy of the EnKF. The accuracy of the data assimilation under the misspecified physical model is also considered. Numerical studies on the Lorenz-96 and the Shallow Water Equation models illustrate that the proposed HD-EnKF algorithms outperform the standard EnKF and widely used inflation methods.
@@ 1 Introduction
Kalman Filter (Kalman, 1960) is a fundamental approach for forecasting and data assimilation for a partially observed dynamic system. Data assimilation is a statistical procedure that combines information from numerical physical model forecasts with observed data to obtain the best possible description of a dynamical system and its uncertainty (Evensen et al., 2022). The assimilated state variables at a time are used to generate the forecast for the next time step. For chaotic systems sensitive to initial values, data assimilation can significantly improve the accuracy of the forecast by using optimized state variables as initial values (Lorenz, 1963). Although much applied in numerical weather and oceanography prediction, Kalman Filter and its later developments (EnKF, three and four-dimensional variational assimilation (3D-Var, 4D-Var; Pailleux, 1990; Courtier et al., 1998)) are the methodological core for scientific algorithms over various fields, including hydrology, remote sensing data inversion, carbon data assimilation, robotic control and automatic driving. Data assimilation methods also provide the underlying algorithms and predictive techniques for building digital twins for social and economic systems.
@@
Estimating the forecast error covariance is a key in the EnKF, which is made in the EnKF by the sample covariance of the forecast ensembles. Past works on the EnKF found that the sampling error can lead to underestimation of the forecast error covariance and can eventually cause filter divergence(Anderson and Anderson, 1999; Constantinescu et al., 2007). Filter divergence refers to a decreasing ability to correct the ensemble state towards the observations after certain rounds of assimilation. One reason for the filter divergence is the variance of the ensemble forecast error becomes too small when there are unaccounted model errors (Lorenc, 2003); the other is due to the substantial mismatch between the estimated and true error covariance matrices (Houtekamer and Mitchell, 1998; Constantinescu et al., 2007).
@@
Along with advances in remote sensing technology in measuring the Earth system, there is a growing availability of high spatio-temporal resolution satellite observations (Stroud et al., 2010; Guinehut et al., 2012; Lellouche et al., 2013). Meanwhile, the high resolution dataset and forecasting are in increasing demands in modern meteorology and atmospheric systems like high resolution weather forecasting, global carbon calculation and ocean data assimilation, which leads to high dimensional geophysical models with high dimensional state vectors. The high dimensional state vectors lead to high dimensional  forecast error covariance and Kalman Gain matrices. The high dimensionality is known to cause the non-convergence of the sample forecast error covariance (Bai and Yin, 1993), felt in the form of filter divergence and unreliable data assimilation in practice. Geophysicists have proposed two methods to counter the problems, the localization method (Furrer et al., 2006; Furrer and Bengtsson, 2007) that sets covariances to zero for the geographical distances bigger than a threshold and the inflation method that enlarges the forecast error covariance by adding either a multiplicative factor or an additive term (Wang and Bishop, 2003; Constantinescu et al., 2007; Liang et al., 2012). However, the statistical properties of these methods are yet to be established. Tong (2018) investigated the performance of the local EnKF with domain localization under linear systems. Katzfuss et al. (2020) proposed Gibbs EnKF and particle EnKF algorithms for high dimensional hierarchical dynamic models from a Bayesian perspective. Al-Ghattas and Sanz-Alonso (2024) developed a non-asymptotic analysis on the difference between the ensembles of single-step Ensemble Kalman Inversion using the true and the localized estimation of forecast error covariances under the framework of effective dimension of matrices and Al-Ghattas et al. (2024) explored the resampling EnKF under linear systems and effective dimension. While Al-Ghattas et al. (2025) studied the difference between the ensembles of single-step EnKF using the true and the thresholded estimation of forecast error covariance functions in a small correlation lengthscale regime. The EnKF methods and their theoretical analysis under high dimensional framework for general non-linear models, especially the theoretical performance of the multiple-step EnKF under both the correct and misspecified models, need further investigation.
This paper proposes several high dimensional EnKF (HD-EnKF) methods and conducts a theoretical analysis of the one-step and multiple-step assimilated state vectors for general nonlinear models and imperfect models under both fixed and high dimensional cases. It is shown that even at the high dimensional setting, the mean square errors (MSEs) of the analysis states by the EnKF could be controlled by the error bounds of the high dimensional forecast error covariance estimators. Thus, consistent recovery of the underlying analysis states can be attained by the proposed HD-EnKF. Specifically, we establish the MSE bounds and weak convergence of the analysis states from the HD-EnKF to the underlying analysis under both fixed and high dimensional cases, as well as for both correct and imperfect physical models. The MSE bounds for the multiple-step forecast and assimilation are also provided. The investigation of the HD-EnKF under the misspecified model gives a more realistic description of the performance of the EnKF. Under the misspecified model, the MSE not only depends on the estimation error of the forecast error covariance matrices but also the discrepancy between the misspecified and the true model. Simulation studies suggest that the proposed HD-EnKF could achieve more accurate and robust assimilation results than the widely used inflation-based EnKF method. This implies that the proposed HD-EnKF is applicable in a range of studies and prospects in earth science like global carbon calculation, weather forecasting and the production of high resolution data sets, considering both the theoretical guarantee and better performances in practice.
@@
The paper is structured as follows. Section 2 reviews the Kalman Filter and the EnKF. Section 3 presents several consistent estimators of the high dimensional forecast error covariance and the key Kalman Gain, and establishes the theoretical properties of the EnKF. The effects of imperfect or misspecified models on the accuracy of the EnKF are provided in Section 4. Results of empirical data assimilation based on two well-known geophysical models are reported in Section 5, followed by a conclusion in Section 6. The proposed HD-EnKF algorithms, additional theoretical and simulation results, and technical details are given in the Supplementary Material (SM).
@@ 2 Review on Kalman and Ensemble Kalman Filters
Assuming $\mathbf{M}_{t}$, $\mathbf{Q}_{t}$, $\mathbf{H}_{t}$, and $\mathbf{R}_{t}$ are known, the Kalman Filter provides a two-step alternative updating framework based on the theory of multivariate normal distribution. In the forecast step, suppose that $\mathbf{x}_{t-1}\mid\mathbf{y}_{1:t-1}\sim{\mathrm{N}}_{p}(\mathbf{x}_{t-1}^{a},\boldsymbol{\Sigma}_{t-1}^{a})$, where $\mathbf{x}_{t-1}^{a}$ is the analysis state vector at the previous time t1t-1 and $\mathbf{y}_{1:t-1}$ denotes the observation $\{\mathbf{y}_{i}\}_{i=1}^{t-1}$ up to time t1t-1, the conditional distribution of $\mathbf{x}_{t}$ given $\mathbf{y}_{1:t-1}$ is
@@
is the forecast error covariance, the variance of the forecast error $\mathbf{x}_{t}-\mathbf{x}_{t}^{f}$.
In the update (assimilation) step, the goal is to update $\mathbf{x}_{t}^{f}$ according to the new observation $\mathbf{y}_{t}$ and to reduce the difference between the true state $\mathbf{x}_{t}$ and the forecast state $\mathbf{x}_{t}^{f}$. The joint distribution of $\left(\mathbf{x}_{t},\mathbf{y}_{t}\right)$ given $\mathbf{y}_{1:t-1}$ is Gaussian, which implies that $\mathbf{x}_{t}\mid\mathbf{y}_{1:t}\sim{\mathrm{N}}_{p}\left({\mathbf{x}}_{t}^{a},{\boldsymbol{\Sigma}}_{t}^{a}\right)$, where, by the conditional Gaussian distribution formulae,
@@
However, models in earth science or engineering are often nonlinear so that $\mathbf{x}_{t}=\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}$ in (2.1), where $\mathcal{M}_{t}$ is a nonlinear model operator that can invalidate the forecast distribution in (2.2). Besides, for large dynamic systems, the dimension of the state vector $\mathbf{x}_{t}$ can be large so that the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ is expensive to compute and store. To tackle these problems, Evensen (1994) proposed the EnKF, where the forecast error covariance is no longer updated as (2.3) but by the sample covariance of $n$ forecast ensembles. It hence only needs to store and operate on $n$ vectors of dimension $p$, and is convenient for parallel computing, and thus improves the computing efficiency and reduces the storage cost.
The EnKF assumes possibly nonlinear model system with a linear observation operator, with the first equation in (2.1) replaced by
$\mathbf{x}_{t}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1}\right)+\mathbf{w}_{t},$ (2.6)
where $\mathcal{M}_{t}$ is the possibly nonlinear model operator.
@@
    $\mathbf{x}_{t,j}^{f}=\mathcal{M}_{t}\left(\mathbf{x}_{t-1,j}^{a}\right)+\mathbf{w}_{t,j},\,\hat{\mathbf{x}}_{t}^{f}=\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{t,j}^{f},\,\mathbf{w}_{t,j}\stackrel{{\scriptstyle IID}}{{\sim}}\mathrm{N}_{p}\left(\mathbf{0},\mathbf{Q}_{t}\right).$ (2.7)
@@
    and obtain the Kalman Gain estimate $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}\big{(}\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t}\big{)}^{-1}$.
@@
    Calculate the perturbed residuals $\mathbf{d}_{t,j}=\mathbf{y}_{t}+\boldsymbol{\varepsilon}_{t,j}-{\mathbf{H}}_{t}\mathbf{x}_{t,j}^{f}$ where $\{\boldsymbol{\varepsilon}_{t,j}\}_{j=1}^{n}\stackrel{{\scriptstyle IID}}{{\sim}}{\mathrm{N}}\left(\mathbf{0},\mathbf{R}_{t}\right)$, and let $\mathbf{d}_{t}=n^{-1}\sum_{j=1}^{n}\mathbf{d}_{t,j}$. Compute the perturbed analysis states and their average, the analysis state,
@@
In practice, the algorithm is initialized at time $t=0$ by drawing $\{\mathbf{x}_{0,j}^{a}\}_{j=1}^{n}$ independently from ${\mathrm{N}}_{p}(\mathbf{x}_{0}^{a},\Sigma_{0}^{a})$ where $\mathbf{x}_{0}^{a}$ and $\Sigma_{0}^{a}$ are either chosen based on the physical model or by pre-running the physical model for a period of time. Then, the ensemble is propagated forward through time, alternating between the forecast and the analysis steps, according to the three-step procedure above. If $\mathbf{R}_{t}$ and $\mathbf{Q}_{t}$ are unknown, their estimates can be entertained to substitute them in (2.7) and (2.9).
The EnKF does not require the expensive recursions (2.3) and (2.4)-(2.5) used in the Kalman Filter. Instead, it estimates the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ by the sample covariance. For the case with non-linear model operator as in (2.6), we can still define the underlying analysis state $\mathbf{x}_{t}^{a}$ by (2.4)-(2.5), where $\mathbf{x}_{t}^{f}=\mathbb{E}\left(\mathbf{x}_{t}\mid\mathbf{y}_{1:t-1}\right)=\mathbb{E}(\mathcal{M}_{t}(\mathbf{x}_{t-1})+\mathbf{w}_{t}\mid\mathbf{y}_{1:t-1})$ is the forecast state and $\boldsymbol{\Sigma}_{t}^{f}$ is the forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}=\operatorname{var}((\mathbf{x}_{t}-\mathbf{x}_{t}^{f})\mid\mathbf{y}_{1:t-1})$. In fact, $\hat{\mathbf{x}}_{t}^{a}$ defined by (2.9) is the solution to the minimizing problem (2.10)
$J(\mathbf{x})=\frac{1}{2}\big{\{}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})^{\mathrm{T}}(\hat{\boldsymbol{\Sigma}}_{t}^{f})^{-1}(\mathbf{x}-\hat{\mathbf{x}}_{t}^{f})+\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)^{\mathrm{T}}\mathbf{R}_{t}^{-1}\left(\mathbf{y}_{t}-\mathbf{H}_{t}\mathbf{x}\right)\big{\}},$ (2.10)
and $\mathbf{x}_{t}^{a}$ defined by (2.4)-(2.5) is the solution to the above minimizing problem with $\hat{\mathbf{x}}_{t}^{f}$ and $\hat{\boldsymbol{\Sigma}}_{t}^{f}$ replaced by $\mathbf{x}_{t}^{f}$ and $\boldsymbol{\Sigma}_{t}^{f}$. For the ease of reference, and to avoid differentiating between the linear and the non-linear cases, we also call $\mathbf{x}_{t}^{a}$ as the analysis state of the Kalman Filter even in the non-linear model case. The next two sections will establish the convergence of the EnKF analysis state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$ of the Kalman Filter.